# Effects of intertube dipole-dipole interactions in nearly integrable one-dimensional 162Dy gases

### v1

@@ II.4 Modeling the effects of the intertube DDI
$U_{\rm DDI}^{\rm inter}(x_{l};y_{l},z_{l})=\sum_{l^{\prime}\neq l}V_{l^{\prime}}^{\rm inter}(x_{l};y_{l},z_{l})\,$, (6)
@@
$V_{l^{\prime}}^{\rm inter}(x_{l};y_{l},z_{l})=\int dx_{l}^{\prime}\frac{\mu_{0}\mu^{2}}{4\pi}\frac{1-3(\hat{r}\cdot\hat{B})^{2}}{r^{3}}n(x_{l^{\prime}};y_{l^{\prime}},z_{l^{\prime}})\,$, (7)
and ${\bf r}=(x_{l}-x_{l^{\prime}},y_{l}-y_{l^{\prime}},z_{l}-z_{l^{\prime}})$.
Specifically, at the decoupling depth $U^{*}_{2D}$ determined in Ref. [42], we use the temperature $T^{*}_{\rm 2D}$ and the density distribution computed in that work to calculate the modification introduced by Eq. (6) to the trapping potentials. We then update the densities so that the system is in thermal equilibrium at the temperature $T^{*}_{\rm 2D}$ in the new trapping potentials. This procedure is then repeated iteratively until the density distribution converges in all the tubes. We therefore end up with density and entropy distributions that differ from the ones in Ref. [42]. Similar iterative procedures, using the new density and entropy distributions, are repeated at each stage at which the temperatures of the 1D gases were determined in the absence of the intertube DDI in Ref. [42]. At the end of our new modeling of the equilibrium state preparation, we have slightly different density and temperature distributions when compared to those found in Ref. [42]. Finally, to model the rapidity measurements, we use generalized hydrodynamics (GHD) [21, 22, 59] to study the expansion of the 1D gases in the weakening effective potential generated by the intertube DDI in the expanding gases.
@@ III Initial state
To analyze in detail the effect of the intertube DDI on the initial state preparation for $\theta_{B}=55^{\circ}$ ($g_{\rm 1D}^{\rm DDI}=0$) and fixed $g_{\rm 1D}^{\rm vdW}$, we split our presentation into two subsections. In Sec. III.1, we discuss the effect of the intertube DDI at the decoupling point and, in Sec. III.2, we discuss the effect of the intertube DDI through the end of the loading process.
III.1 1D gases at the decoupling point $U^{*}_{2D}$
@@
In our calculations, we use the optimal values $U^{*}_{2D}=5E_{R}$ and $T^{*}_{2D}=25$ nK determined in Ref. [42]. Note that $T^{*}_{2D}=25$ nK is lower than $T_{3D}\sim 38$ nK, i.e., there is cooling when one transitions between three dimensions and one dimension [42]; see also Ref. [60].
@@
In Figs. 2(a) and 2(b), we show the atom distribution across all the tubes (dots) without and with the intertube DDI correction. The color indicates the number of atoms $N_{l}$ in each tube (see color scale on the right). The distributions are nearly indistinguishable from each other. In Fig. 2(c), we plot the integrated atom distributions along the $y$ (blue lines) and $z$ (green lines) directions, which are also nearly indistinguishable from each other. The asymmetry in the size of the cloud in the different directions reflects the asymmetry of the employed trapping potential.
@@
Figure 2: Atom distribution at decoupling. (a) [(b)] Distribution of atoms $N_{\text{atom}}(y_{l},z_{l})$ in the tubes when we neglect (account for) the intertube DDI. Each dot indicates a tube “ll” at position $(y_{l},z_{l})$. (c) Distribution of atoms along the $y$ direction, $N_{y}=\sum_{z_{l}}N_{\text{atom}}(y_{l},z_{l})$ (blue lines), and along the $z$ direction, $N_{z}=\sum_{y_{l}}N_{\text{atom}}(y_{l},z_{l})$ (green lines), in calculations with (dashed) and without (solid) the intertube DDI correction. (d) The number of tubes $N_{\text{tube}}$ with $N_{\text{atom}}$ atoms versus $N_{\text{atom}}$.
@@
To identify the nature of the differences between modeling with and without the intertube DDI, in Fig. 2(d) we plot the number of tubes $N_{\rm tube}$ whose number of atoms is $N_{\rm atom}$ as a function of $N_{\rm atom}$. Since our modeling is done within the LDA, $N_{l}$ can take noninteger values in the tubes. To make the histogram shown in Fig. 2(d), $N_{l}$ is rounded to the closest integer ($N_{\rm tube}$ for $N_{\rm atom}=0$ in the plot is the number of tubes with $0<N_{l}<0.5$). Figure 2(d) shows that the intertube DDI decreases the maximal occupation of the tubes, i.e., in our asymmetric geometry [see Figs. 2(a)–2(c)] it acts like a weak antitrap in the $y$-$z$ plane.
@@
Figure 3: Density and rapidity distributions at decoupling. (a) Average density distribution of the 1D gases with (dashed lines) and without (solid lines) including the intertube DDI correction at decoupling. The green solid line in the top panel shows the difference $\Delta n(x)=n(x;{\rm with\,}U^{\rm intra}_{\rm DDI})-n(x;{\rm without\,}U^{\rm intra}_{\rm DDI})$. (b) Same as (a) but for rapidity distribution $f(\theta)$, with $\Delta f(\theta)=f(\theta;{\rm with\,}U^{\rm intra}_{\rm DDI})-f(\theta;{\rm without\,}U^{\rm intra}_{\rm DDI})$.
@@
The main observables of interest in this work, the averaged density $n(x)$ and the rapidity distribution $f(\theta)$, are shown at the decoupling point in the main panels of Figs. 3(a) and 3(b), respectively. The differences between the results without and with the intertube DDI are small. To highlight their nature, we magnify them at the top of each panel in Fig. 3. They show that, in the presence of the intertube DDI, the density is slightly lower about the center of the trap (the density distribution is therefore slightly broader), and the occupation of the low rapidity modes is slightly higher (the rapidity distribution is slightly narrower). Hence, like in the $y$-$z$ plane, the intertube DDI in our asymmetric geometry acts like a weak antitrap along the $x$ direction.
@@
A noticeable feature of Fig. 2(d) obtained from finite-temperature modeling within the LDA is the large number of tubes containing less than two atoms. Despite their abundance, they only account for less than $5\%$ of the total number of atoms in the system. Given the self-consistent nature of the calculations done in this work to account for the intertube DDI, including those tubes is computationally very demanding and they almost do not change the results: Their density is very low and they are located at the edge of the system, and thus, they contribute very little to the total intertube DDI energy. Also, ignoring those tubes almost does not change the average rapidity distribution in the absence of the intertube DDI (see Appendix A). Therefore, we ignore them in the remainder of this paper, both in the calculations in which the intertube DDI is included as well as those in which it is not. We stress that, in contrast to the modeling in Ref. [42], in this work we do not round $N_{l}$ to the closest integer in any of our calculations. In Appendix A we show that, in the absence of the intertube DDI, the rapidity distribution obtained with no rounding is closer to the experimental one than the one obtained in Ref. [42] with rounding, so here we remove this additional simplification of Ref. [42].
@@
After the tubes decouple, we assume that the further increase in the depth of the 2D optical lattice, the canceling of the antitrapping potential, and the reduction of the trapping frequency along the $x$ direction ($f_{x}$: $55.5$ $\to 36.4 Hz$) take place adiabatically. That is, we compute the final temperature $T_{l}$ of each tube at the end of the loading assuming that the entropy of each tube is the same as that at the decoupling point $S_{l}$. In practice, we search for $T_{l}$ (on a temperature grid with 0.5 nK steps) such that the entropy matches $S_{l}$ at the final $g_{\rm 1D}$ and 1D trapping potential including the intertube DDI modification. The intertube DDI correction introduced at the decoupling point affects this step through the modification of $N_{l}$ and $S_{l}$. In addition, the intertube DDI modifies the trapping potential, requiring an iterative calculation of $n_{l}(x)$ and $T_{l}$ until convergence is achieved. The effective contact interaction at the end of state preparation is $g^{\rm vdW}_{\rm 1D}=8.5$ $\frac{\hbar^{2}}{\text{m}}\,\mu\text{m}^{-1}$, and we find that the corresponding average effective dimensionless parameter to be $\gamma_{T}=6.7$ (6.5) in the presence (absence) of the intertube DDI. This is nearly unchanged from $\gamma_{T}=6.7$ obtained in Ref. [42]. A motivation and detailed discussion of our definition of $\gamma_{T}$ in 2D arrays of inhomogeneous finite-temperature 1D gases can be found Ref. [42].
@@
Figure 4: Temperature distribution at the end of the loading process. (a) [(b)] Temperature for tubes without (with) intertube DDI after the state preparation. Each dot indicates a 1D tube at position $(y,z)$. (c) [(d)] The ratio between atoms at temperature $T$ ($N_{T}$) and the total number of atom ($N_{\text{total}}$), after the state preparation without (with) intertube DDI correction. The temperature is binned in $0.5$ nK intervals. We show results for tubes with $N_{l}\geq 2$ atoms.
@@
Figures 4(a) and 4(b) show the temperature $T_{l}$ distribution without and with intertube DDI correction, respectively, for all tubes with $N_{l}\geq 2$. Figures 4(c) and 4(d) show the corresponding distributions of the fraction of atoms $N_{T}/N_{\rm tot}$ with a temperature $T$ plotted as functions of $T$. Due to the reduction of trapping potential in the $x$ direction in the final stage of state preparation, the temperature of the tubes decreases significantly from the value at decoupling ($T^{*}=25$ nK) in both cases. In this final step of the loading process, we find that the intertube DDI correction has the clear effect of lowering the temperature of the tubes, specially of those that have lower filling (the ones away from the center of the system). This is again consistent with the intertube DDI acting as an antitrap along $x$, as we found at the decoupling point.
Figure 5: Density and rapidity distributions at the end of the loading process. (a) Density distributions $n(x)$ at the end of the loading process without (red dotted line) and with (blue dashed dotted line) the intertube DDI correction. (b) Rapidity distributions $f(\theta)$ at the end of the loading process without (red dotted line) and with (blue dashed dotted line) the intertube DDI correction. We also show the rapidity distribution obtained after a 15 ms expansion in the tubes (green dashed line, overlapping with the red dotted line), and the experimental results (orange solid line) from Ref. [42].
@@
In Fig. 5(a), we show the average density distributions $n(x)$ at the end of the state preparation without (red dotted line) and with (blue dashed dotted line) the intertube DDI correction. They are very close to each other despite the clearly different temperatures distributions seen in Fig. 4. This can be understood by considering that the lower temperature decreases the width of density distribution while the effective antitrapping potential generated by the intertube DDI increases it. Those two effects compete with each other, resulting in a density distribution in the presence of the DDI that is very close to that in the absence of the DDI. In this particular case, the aforementioned competition results in a density distribution that is slightly narrower in the presence of the DDI than in its absence.
@@
In Fig. 5(b), we show the corresponding average rapidity distributions. For the rapidity distribution, the intertube DDI correction to the temperatures (lowering the temperature) and the trapping potentials (antitrapping) both lower the width of the rapidity distribution. Hence, we find that the narrowing of the rapidity distribution due to the intertube DDI is more significant than that of the density distribution. In Fig. 5(b), we also show the rapidity distribution measured in Ref. [42]. One can see that if the effect of the intertube DDI is accounted for only in the state preparation (as we have done so far here), the modeled rapidity distribution (slightly) differs more from the experimental one than the one obtained in the absence of the DDI. As mentioned in Ref. [42], one also needs to account for the effect of the intertube DDI during the expansion, which we do next.
@@ IV Rapidity measurement
As mentioned in Sec. II.1, to measure the rapidity distribution the 1D trapping potential is suddenly turned off ($V_{\rm 1D}\to 0$) and the 1D gases are allowed to expand for a time $t_{\rm ev}$ while the 2D lattice $U_{\rm 2D}$ is still on. During the expansion, the atoms interact and their momentum distribution changes with time, asymptotically approaching the rapidity distribution at long times [40, 41, 49, 50, 52, 53]. In the absence of a DDI, i.e., for atoms interacting via only contact interactions, the rapidity distribution is conserved during the expansion. Therefore, measuring the momentum distribution after a long $t_{\rm ev}$ is equivalent to measuring the rapidity distribution of the initial state. In the presence of the DDI, the rapidity distribution changes during the expansion. Since the intertube DDI acts like an effective antitrapping potential, it accelerates the quasiparticles and broadens the rapidity distribution. Namely, the effect of the intertube DDI on the rapidity distribution during the expansion dynamics is the opposite to the effect it has in the initial state.
We use GHD [21, 22, 59] to compute the density $n(x;t_{\rm ev})$ and the rapidity distribution $f(\theta;t_{\rm ev})$ at $t_{\rm ev}=15$ ms, which is the 1D expansion time in the experiments before the 2D lattice is turned off and the momentum distribution of the atoms (which we assume is the same as the rapidity distribution at that time) is measured using a standard TOF expansion. In our simulation, all the 1D gases expand in a weakening anti-trap generated by the DDI [see Eq. (6)], which we update throughout the expansion.
In Fig. 6(a), we show the density distribution $n(x)$ at different times $t_{\rm ev}$ during the expansion in 1D in the presence of the intertube DDI. After the 1D trap is turned off, the atom cloud expands ballistically and by $t_{\rm ev}=15$ ms it has a size that is about $4$ times the initial size, as needed for the momentum distribution to evolve into the rapidity distribution [40]. By that time, we find that the remaining intertube DDI energy is about 1.5% of the energy in the initial state. Namely, almost all of the intertube DDI energy has been converted into kinetic energy. Figure 6(b) shows the corresponding evolution of the rapidity distribution. The changes in $f(\theta)$ occur mostly within expansion times $t_{\rm ev}\lesssim 10$ ms. We find the rapidity distribution to be nearly stationary by $t_{\rm ev}=15$ ms and, as expected, broader than the initial rapidity distribution.
Figure 6: Density and rapidity distributions during the expansion. (a) $f(x)$ at 1D expansion time $t_{\rm ev}$. We show results for $t_{\rm ev}=$ 0 ms (blue dashed dotted line), 5 ms (black dashed line), 10 ms (red dotted line), and 15 ms (green solid line). (b) Same as (a) but for $f(\theta)$.

### v2

@@ II.4 Modeling the effects of the intertube DDI
$U_{\rm DDI}^{\rm inter}(x_{\ell};y_{\ell},z_{\ell})=\sum_{\ell^{\prime}\neq\ell}V_{\ell^{\prime}}^{\rm inter}(x_{\ell};y_{\ell},z_{\ell})\,$, (6)
@@
$V_{\ell^{\prime}}^{\rm inter}(x_{\ell};y_{\ell},z_{\ell})=\int dx_{\ell^{\prime}}\frac{\mu_{0}\mu^{2}}{4\pi}\frac{1-3(\hat{r}\cdot\hat{B})^{2}}{r^{3}}n(x_{\ell^{\prime}};y_{\ell^{\prime}},z_{\ell^{\prime}})\,$, (7)
[…]
Specifically, at the decoupling depth $U^{*}_{\rm 2D}$ determined in Ref. [43], we use the temperature $T^{*}_{\rm 2D}$ and the density distribution computed in that work to calculate the modification introduced by Eq. (6) to the trapping potentials. We then update the densities so that the system is in thermal equilibrium at the temperature $T^{*}_{\rm 2D}$ in the new trapping potentials. This procedure is repeated iteratively until the density distribution converges in all the tubes. We therefore end up with density and entropy distributions that differ from the ones in Ref. [43]. Similar iterative procedures, using the new density and entropy distributions, are repeated at each stage at which the temperatures of the 1D gases were determined in the absence of the intertube DDI in Ref. [43]. At the end of our new modeling of the equilibrium state preparation, we have slightly different density and temperature distributions when compared to those found in Ref. [43]. Finally, to model the rapidity measurements, we use GHD [21, 22, 58] to study the expansion of the 1D gases in the weakening effective potential generated by the intertube DDI in the expanding gases.
@@ III Initial state
To analyze in detail the effect of the intertube DDI on the initial state preparation for $\theta_{B}=55^{\circ}$ ($g_{\rm 1D}^{\rm DDI}=0$) and fixed $g_{\rm 1D}^{\rm vdW}$, we split our presentation into two sections. In Sec. III.1 we discuss the effect of the intertube DDI at the decoupling point and in Sec. III.2 we discuss the effect of the intertube DDI through the end of the loading process.
III.1 1D gases at the decoupling point $U^{*}_{\rm 2D}$
@@
In our calculations, we use the optimal values $U^{*}_{\rm 2D}=5E_{R}$ and $T^{*}_{\rm 2D}=25$ nK determined in Ref. [43]. Note that $T^{*}_{\rm 2D}=25$ nK is lower than $T_{3D}\sim 38$ nK, i.e., there is cooling when one transitions between three dimensions and one dimension [43] (see also Ref. [59]).
@@
In Figs. 2(a) and 2(b) we show the atom distribution across all the tubes (dots) without and with the intertube DDI correction. The color indicates the number of atoms $N_{\ell}$ in each tube (see color scale on the right). The distributions are nearly indistinguishable from each other. In Fig. 2(c) we plot the integrated atom distributions in the $y$ (blue lines) and $z$ (green lines) directions, which are also nearly indistinguishable from each other. The asymmetry in the size of the cloud in the different directions reflects the asymmetry of the employed trapping potential.
@@
Figure 2: Atom distribution at decoupling. Distribution of atoms $N_{\text{atom}}(y_{\ell},z_{\ell})$ in the tubes when we (a) neglect and (b) account for the intertube DDI. Each dot indicates a tube $\ell$ at position $(y_{\ell},z_{\ell})$. (c) Distribution of atoms in the $y$ direction, $N_{y}=\sum_{z_{\ell}}N_{\text{atom}}(y_{\ell},z_{\ell})$ (blue lines), and in the $z$ direction, $N_{z}=\sum_{y_{\ell}}N_{\text{atom}}(y_{\ell},z_{\ell})$ (green lines), in calculations with (dashed) and without (solid) the intertube DDI correction. (d) Number of tubes $N_{\text{tube}}$ with $N_{\text{atom}}$ atoms vs $N_{\text{atom}}$.
@@
To identify the nature of the differences between modeling with and without the intertube DDI, in Fig. 2(d) we plot the number of tubes $N_{\rm tube}$ whose number of atoms is $N_{\rm atom}$ as a function of $N_{\rm atom}$. Since our modeling is done within the LDA, $N_{\ell}$ can take noninteger values in the tubes. To make the histogram shown in Fig. 2(d), $N_{\ell}$ is rounded to the closest integer ($N_{\rm tube}$ for $N_{\rm atom}=0$ in the plot is the number of tubes with $0<N_{\ell}<0.5$). Figure 2(d) shows that the intertube DDI decreases the maximal occupation of the tubes, i.e., in our asymmetric geometry [see Figs. 2(a)–2(c)] it acts like a weak antitrap in the $y$-$z$ plane.
@@
Figure 3: Density and rapidity distributions at decoupling. (a) Average density distribution of the 1D gases with (dashed lines) and without (solid lines) including the intertube DDI correction at decoupling. The green solid line in the top panel shows the difference $\Delta n(x)=n(x;{\rm with\,}U^{\rm inter}_{\rm DDI})-n(x;{\rm without\,}U^{\rm inter}_{\rm DDI})$. (b) Same as (a) but for rapidity distribution $f(\theta)$, with $\Delta f(\theta)=f(\theta;{\rm with\,}U^{\rm inter}_{\rm DDI})-f(\theta;{\rm without\,}U^{\rm inter}_{\rm DDI})$.
@@
The main observables of interest in this work, the averaged density $n(x)$ and the rapidity distribution $f(\theta)$, are shown at the decoupling point in the main panels of Figs. 3(a) and 3(b), respectively. The differences between the results without and with the intertube DDI are small. To highlight their nature, we magnify them at the top of each panel in Fig. 3. They show that, in the presence of the intertube DDI, the density is slightly lower about the center of the trap (the density distribution is therefore slightly broader) and the occupation of the low rapidity modes is slightly higher (the rapidity distribution is slightly narrower). Hence, like in the $y$-$z$ plane, the intertube DDI in our asymmetric geometry acts like a weak antitrap in the $x$ direction.
@@
A noticeable feature of Fig. 2(d) obtained from finite-temperature modeling within the LDA is the large number of tubes containing less than two atoms. Despite their abundance, they only account for less than $5\%$ of the total number of atoms. Given the self-consistent nature of the calculations done in this work to account for the intertube DDI, including those tubes is computationally very demanding and they almost do not change the results: Their density is very low and they are located at the edge of the system; thus, they contribute very little to the total intertube DDI energy. Also, ignoring those tubes almost does not change the average rapidity distribution in the absence of the intertube DDI (see the Appendix). Therefore, we ignore them in the remainder of this paper, in both the calculations in which the intertube DDI is included and those in which it is not. We stress that, in contrast to the modeling in Ref. [43], in this work we do not round $N_{\ell}$ to the closest integer in any of our calculations. In the Appendix we show that, in the absence of the intertube DDI, the rapidity distribution obtained with no rounding is closer to the experimental one than the one obtained in Ref. [43] with rounding, so here we remove this additional simplification of Ref. [43].
@@
After the tubes decouple, we assume that the further increase in the depth of the 2D optical lattice, the canceling of the antitrapping potential, and the reduction of the trapping frequency in the $x$ direction ($f_{x}$: $55.5$ $\to 36.4 Hz$) take place adiabatically for each tube independently. (The global thermalizing effects of the intertube interactions occur in a much longer time scale [15].) More specifically, we compute the final temperature $T_{\ell}$ of each tube at the end of the loading assuming that the entropy of each tube is the same as that at the decoupling point $S_{\ell}$. In practice, we search for $T_{\ell}$ (on a temperature grid with 0.5 nK steps) such that the entropy matches $S_{\ell}$ at the final $g_{\rm 1D}$ and 1D trapping potential including the intertube DDI modification. The intertube DDI correction introduced at the decoupling point affects this step through the modification of $N_{\ell}$ and $S_{\ell}$. In addition, the intertube DDI modifies the trapping potential, requiring an iterative calculation of $n_{\ell}(x)$ and $T_{\ell}$ until convergence is achieved. The effective contact interaction at the end of state preparation is $g^{\rm vdW}_{\rm 1D}=8.5$ $\frac{\hbar^{2}}{m}\,\mu\text{m}^{-1}$, and we find that the corresponding average effective dimensionless parameter is $\gamma_{T}=6.7$ (6.5) in the presence (absence) of the intertube DDI. This is nearly unchanged from $\gamma_{T}=6.7$ obtained in Ref. [43]. A motivation and detailed discussion of our definition of $\gamma_{T}$ in 2D arrays of inhomogeneous finite-temperature 1D gases can be found Ref. [43].
@@
Figure 4: Temperature distribution at the end of the loading process. Temperature for tubes (a) without and (b) with intertube DDI after the state preparation. Each dot indicates a 1D tube at position $(y,z)$. Also shown is the ratio between atoms at temperature $T$ ($N_{T}$) and the total number of atom $N_{\rm tot}$ after the state preparation (c) without and (d) with intertube DDI correction. The temperature is binned in $0.5$-nK intervals. We show results for tubes with $N_{\ell}\geq 2$ atoms.
@@
Figures 4(a) and 4(b) show the temperature $T_{\ell}$ distribution without and with intertube DDI correction, respectively, for all tubes with $N_{\ell}\geq 2$. Figures 4(c) and 4(d) show the corresponding distributions of the fraction of atoms $N_{T}/N_{\rm tot}$ with a temperature $T$ plotted as a function of $T$. Due to the reduction of trapping potential in the $x$ direction in the final stage of state preparation, the temperature of the tubes decreases significantly from the value at decoupling ($T^{*}=25$ nK) in both cases. In this final step of the loading process, we find that the intertube DDI correction has the clear effect of lowering the temperature of the tubes, specially of those that have lower filling (the ones away from the center of the system). This is again consistent with the intertube DDI acting as an antitrap in the $x$ direction, as we found at the decoupling point.
@@
In Fig. 5(a) we show the average density distributions $n(x)$ at the end of the state preparation without (red dotted line) and with (blue dashed dotted line) the intertube DDI correction. They are very close to each other despite the clearly different temperature distributions seen in Fig. 4. This can be understood by considering that the lower temperature decreases the width of the density distribution while the effective antitrapping potential generated by the intertube DDI increases it. Those two effects compete with each other, resulting in a density distribution in the presence of the DDI that is very close to that in the absence of the DDI. In this particular case, the aforementioned competition results in a density distribution that is slightly narrower in the presence of the DDI than in its absence.
Figure 5: Density and rapidity distributions at the end of the loading process. (a) Density distributions $n(x)$ at the end of the loading process without (red dotted line) and with (blue dash-dotted line) the intertube DDI correction. (b) Rapidity distributions $f(\theta)$ at the end of the loading process without (red dotted line) and with (blue dash-dotted line) the intertube DDI correction. We also show the rapidity distribution obtained after a 15 ms expansion in the tubes (green dashed line, overlapping with the red dotted line), and the experimental results (orange solid line) from Ref. [43].
@@
In Fig. 5(b) we show the corresponding average rapidity distributions. For the rapidity distribution, the intertube DDI correction to the temperatures (lowering the temperature) and the trapping potentials (antitrapping) both lower the width of the rapidity distribution. Hence, we find that the narrowing of the rapidity distribution due to the intertube DDI is more significant than that of the density distribution. In Fig. 5(b) we also show the rapidity distribution measured in Ref. [43]. One can see that when the effect of the intertube DDI is included only in the state preparation (as we have done so far), the modeled rapidity distribution (slightly) differs more from the experimental one than that obtained in the absence of the DDI. As mentioned in Ref. [43], one also needs to include the effect of the intertube DDI during the expansion, which we do next.
@@ IV Rapidity measurement
To measure the rapidity distribution the 1D trapping potential is suddenly turned off ($V_{\rm 1D}\to 0$) and the 1D gases are allowed to expand for a time $t_{\rm ev}$ while the 2D lattice $U_{\rm 2D}$ is still on. During the expansion, the atoms interact and their momentum distribution changes with time, asymptotically approaching the rapidity distribution at long times [41, 42, 49, 50, 52, 53]. In the absence of a DDI, i.e., for atoms interacting only via contact interactions, the rapidity distribution is conserved during the expansion. Therefore, measuring the momentum distribution after a long $t_{\rm ev}$ is equivalent to measuring the rapidity distribution of the initial state. In the presence of the DDI, the rapidity distribution changes during the expansion. Since the intertube DDI acts like an effective antitrapping potential, it accelerates the quasiparticles and broadens the rapidity distribution, namely, the effect of the intertube DDI on the rapidity distribution during the expansion dynamics is the opposite of the effect it has in the initial state.
We use GHD [21, 22, 58] to compute the density $n(x;t_{\rm ev})$ and the rapidity distribution $f(\theta;t_{\rm ev})$ at $t_{\rm ev}=15$ ms, which is the 1D expansion time in the experiments before the 2D lattice is turned off and the momentum distribution of the atoms (which we assume is the same as the rapidity distribution at that time) is measured using a standard TOF expansion. In our simulation, all the 1D gases expand in a weakening antitrap generated by the DDI [see Eq. (6)], which we update throughout the expansion.
In Fig. 6(a) we show the density distribution $n(x)$ at different times $t_{\rm ev}$ during the expansion in one dimension in the presence of the intertube DDI. After the 1D trap is turned off, the atom cloud expands ballistically and by $t_{\rm ev}=15$ ms it has a size that is about four times the initial size, as needed for the momentum distribution to evolve into the rapidity distribution [41]. By that time, we find that the remaining intertube DDI energy is about 1.5% of the energy in the initial state, namely, almost all of the intertube DDI energy has been converted into kinetic energy. Figure 6(b) shows the corresponding evolution of the rapidity distribution. The changes in $f(\theta)$ occur mostly within expansion times $t_{\rm ev}\lesssim 10$ ms. We find the rapidity distribution to be nearly stationary by $t_{\rm ev}=15$ ms and, as expected, broader than the initial rapidity distribution.
Figure 6: Density and rapidity distributions during the expansion. (a) Density distribution $n(x)$ at 1D expansion time $t_{\rm ev}$. We show results for $t_{\rm ev}=$ 0 ms (blue dash-dotted line), 5 ms (black dashed line), 10 ms (red dotted line), and 15 ms (green solid line). (b) Same as (a) but for the rapidity distribution $f(\theta)$.